OpenStudy (cutiecomittee123):

what is the maximum value of y=-3+4cos(5pi/6(x+4))

jimthompson5910 (jim_thompson5910):

cos(x) maxes out at 1. It's the largest it can get we can replace x with any complicated expression we want so cos(BLAH) also maxes out at 1 What does this mean? It means that if we want to max out y, we just replace `cos(5pi/6(x+4))` with the largest numeric value it can be, which is 1

jimthompson5910 (jim_thompson5910):

so `y=-3+4cos(5pi/6(x+4))` turns into `y=-3+4(1)` from here you just calculate `-3+4(1)`

OpenStudy (cutiecomittee123):

so -12

OpenStudy (cutiecomittee123):

I mean -1

OpenStudy (cutiecomittee123):

oh wait lol brain fart 1

jimthompson5910 (jim_thompson5910):

yes, in the end, the max is y = 1

OpenStudy (cutiecomittee123):

sweet thanks:)

jimthompson5910 (jim_thompson5910):

this graph confirms the answer. Notice how the peak of each mountain is at y = 1 https://www.desmos.com/calculator/lboel2wxy6

OpenStudy (cutiecomittee123):

I see it, good to know thanks for posting that because I am a visual learner and that helped my understanding

jimthompson5910 (jim_thompson5910):

glad it helps out